OpenStudy (anonymous):

find the x and y intercepts of f(x)=-9+3(x+1)+^2

OpenStudy (anonymous):

What is +^2?

OpenStudy (anonymous):

can you rewrite the original problem?

OpenStudy (anonymous):

To find x - intercept let y = 0 and solve. To find y - intercept let x = 0 and solve.

OpenStudy (anonymous):

f (x)=-9+3(x+1)

OpenStudy (anonymous):

0 = 9+3x+3 f(x) = 9+3(0)+3

OpenStudy (anonymous):

Solve

OpenStudy (anonymous):

how can I find y if there is no y

OpenStudy (anonymous):

f(x) = y

OpenStudy (anonymous):

im little confuse about finding y.i can find x

OpenStudy (anonymous):

f(x) = y, like I said, f(x) is y, just change it to y.

OpenStudy (anonymous):

y = -9+3(x+1)

OpenStudy (anonymous):

ok thanks for ur help

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

Describe the end behavior of f(x)=-3^2+5x^3-7x+1